OpenStudy (anonymous):

prove : log (n+1) - log n > 3/(10n)

OpenStudy (anonymous):

Use this on LHS: \[\large Log(a) - Log(b) = Log(\frac{a}{b})\]

ganeshie8 (ganeshie8):

looks tough

OpenStudy (anonymous):

y i see.. but i stuck in next step :(

ganeshie8 (ganeshie8):

is that a prove : or solve ?

OpenStudy (anonymous):

prove, not solve

OpenStudy (anonymous):

may be you can try induction method

OpenStudy (anonymous):

how to use it??

OpenStudy (anonymous):

yu know the induction principle used in+1????

OpenStudy (anonymous):

when the formula is true ,you should use the method of mathematical proof,assuming that formula is true when 'n'.then prove it is true when 'n+1'.

OpenStudy (anonymous):

firstly proving the result is true for n=1 then assuming it true for n=k,any positive integer then to prove it true for n=k+1 using the result for n=k

OpenStudy (anonymous):

for n=1, log 2 > 3/10 0.301... > 0.3 (right) for n=k, => log((k+1)/k)) > 3/(10k) (assuming it true) for n=k+1, => log((k+2)/(k+1)) > 3/(10k+10) what is the next step??

OpenStudy (anonymous):

i told you using kth one,we have to prove for n=k+1 so replace k with k+1 in log(k+1/k)>3/10k,,,,,,,,,and the result is there

OpenStudy (anonymous):

since it is true for any positive integer k+1 ,,,,,,,,,it is true for all n>0

OpenStudy (anonymous):

got it???????

OpenStudy (anonymous):

sorii i cant prove for n=k+1, help me please..

OpenStudy (anonymous):

n is positive integer?

OpenStudy (anonymous):

not rule that n must be positive integer... i think from log (n+1), then n+1>0 or n>-1, maybe...